OpenStudy (anonymous):

Solve the compound inequality 2x-1/10≤-7 or 2x-1/10≥4

OpenStudy (amistre64):

in equalities are best solved with graphing so you can see the area of the solutions

OpenStudy (amistre64):

2x-1 2x-1 ----- = -7 or ------ = 4 10 10 graph these two curves as boundaries

OpenStudy (anonymous):

but I don't know how to solve it.

OpenStudy (anonymous):

I don't have to graph them for this problem only solve.

OpenStudy (amistre64):

is that (1/10) or is the 10 under it all like a denom?

OpenStudy (anonymous):

10 under like a denom

OpenStudy (amistre64):

it still shows up the same on wolfram

OpenStudy (anonymous):

(2x-1)/10≤-7 or (2x-1)/10≥4

OpenStudy (amistre64):

and I assumed it was a y in there someplce; but this is really just on the number line right :)

OpenStudy (anonymous):

i saw the graph but i don't understand what I am looking for.

OpenStudy (anonymous):

there is no Y and I don't have to graph anything I just have to solve the compound inequality.

myininaya (myininaya):

\[\frac{2x-1}{10} \le -7\] \[2x-1 \le -70\] \[2x \le -69\] \[x \le \frac{-69}{2}\] -------------------- \[\frac{2x-1}{10} \ge 4\] \[2x-1 \ge 40\] \[2x \ge 41\] \[x \ge \frac{41}{2}\] ~~~~~~* *~~~~~~ ________|_______|________ -69/2 41/2 x is an element of (-inf,-69/2]U[41/2,inf)

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

\begin{array} \frac{2x-1}{10}≤-7&&\frac{2x-1}{10}≥4&; *10\\ &&\\ 2x-1 \le-70&&2x-1\ge40&;+1\\ &&\\ 2x\le-69&&2x\ge41&;/2\\ &&\\ x\le\frac{-69}{2}&&x\ge\frac{41}{2} \end{array} \begin{align} (<------69/2]........[41/2 ------>) \end{align}

OpenStudy (amistre64):

lol .... all that work and I typoed the first line

OpenStudy (anonymous):

it's ok. thank you for your help!

OpenStudy (anonymous):

do you think you can help me with a work problem I have been struggling with?

OpenStudy (amistre64):

youre welcome :) the answer should appear in your matterial as: \[\left(-\infty,-\frac{69}{2}\right]\ U\ \left[\frac{41}{2},\infty\right)\]

OpenStudy (amistre64):

sure can maybe

OpenStudy (anonymous):

I Think it boggles a lot of people because I have been trying to get help on it for days!

OpenStudy (amistre64):

im used to being confused, so let me see it :)

OpenStudy (anonymous):

Bayside insurance offers two health plans. Under plan A, Giselle would have to pay the first $80 of her medical bills plus 40% of the balance. Plan B Giselle would pay first $210 but only 30% of the remaining balance. For what amount of medical bills will plan B save Giselle Money? Assume she has over $210 in bills

OpenStudy (amistre64):

this ones been around here before... its tricky if you over look some facts

OpenStudy (anonymous):

yea that's because I have been posting it and I can't get anyone to help me! lol

OpenStudy (amistre64):

plan A: Cost = 80+(Bill-$80)(.40) Plan B: Cost = 120+(Bill-$210)(.30) Do you see why its tricky ?

OpenStudy (amistre64):

the 120 sposed to be 210 lol

OpenStudy (anonymous):

ahhh

OpenStudy (amistre64):

the equations read: Cost to G = 80 plus whatevers left if the bill times 40 percent

OpenStudy (anonymous):

right

OpenStudy (amistre64):

lets clean these up some: plan A = 80+Bill(.4)-80(.4) = Bill(.4) +48 Plan B = 210+Bill(.3)-210(.3) = Bill(.3) + 147 when these 2 plans match, they flip over each other :)

OpenStudy (amistre64):

Bill(.3) + 147 = Bill(.4) + 48 147 - 48 = Bill(.4) - Bill(.3) 99 = Bill(.1) 99/.1 = Bill = 990 ; iff I did it righrt

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3Dx%28.4%29%2B48+and+y%3Dx%28.3%29%2B147 this shows where they cross at and says that x = 990 as well

OpenStudy (amistre64):

at a bill of 990; the cost of each plan to G is abt 444

OpenStudy (anonymous):

so with plan B that is the amount of money she will save .

OpenStudy (anonymous):

just checked and that is right....THANK YOU!!!!! You deserve a few more medals. I owe you some! I have been trying that one for days!!!! I am so releaved!